HETAV PATEL

Grade 11,

Integral sec^2(x)/cosec^2(x)

Integral sec^2(x)/cosec^2(x)

Grade:Upto college level

4 Answers

Rathod Shankar AskiitiansExpert-IITB
13 years ago

dear anitha

 

it will become integral{(tan(x))^2}dx

integral   {(tan(x))^2}dx


= integral   {(sec(x))^2 - 1}dx       [since tan^2(x)+1=sec^2(x)]


= 1/2( tan(x)) -1        [integral{sec^2(x)=tanx]

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Rathod Shankar

IIT Bombay

AKASH GARG
13 years ago

tan^2x=sec^2x-1

integrate it,

=> tanx-x                   ans.

subhangee sahoo
12 years ago

integration sec^2(x)/cosec^2(x)

= intg 1/cos^2(x)/1/sin^2(x)

= intg tan^2(x)

=intg sec^2(x)- intg dx

=tanx-x+c

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

it will becomeintegral{(tan(x))^2}dx
integral {(tan(x))^2}dx
= integral {(sec(x))^2 - 1}dx [since tan^2(x)+1=sec^2(x)]
= 1/2( tan(x)) – X +C [integral{sec^2(x)=tanx]

Thanks and Regards